OpenStudy (anonymous):

(5(1/5))/(6(1/6))=t/(7(1/7))

OpenStudy (anonymous):

Is The 5,6 And 7 A Whole Number?

OpenStudy (anonymous):

\[5\frac{ 1 }{ 5} \div \frac{ 1 }{ 6}= t \div 7\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

yes!